Suppose that a change ? ( 2 ) in the price of bread from R10 to R12.50 leads to a change in the quantity demanded of bread from 65 90 units. Use the arc units to elasticity method to calculate and interpret the price elasticity of demand for bread ( 4 )
1
Expert's answer
2020-10-09T07:30:17-0400

PEd ="-0.38462\over 0.25" =-1.5384 which is different from 0.25


Arc Ed="\dfrac{[\dfrac{(Qd\scriptstyle 2-Qd\scriptstyle 1)}{midpoint~Qd}]}{[\tfrac{P_2-P_1}{midpoint P}]}"


=midpoint Qd ="{Qd_1+Qd_2 \over 2}" ="{65+90 \over 2}" = 77.5


midpoint price ="{P_1+P_2 \over 2}" ="{12.5+10 \over 2}" =11.25


% change in quantity demanded ="(90-65) \over 77.5" =0.3226

"\% ~change ~in ~price" ="(10-12.5)\over 11.25" =-0.22222

Arc Ed ="\dfrac{0.3226}{-0.2222}"

=-1.4517

This shows that increase in price by 1 unit reduces demand unit by 1.4517